OpenStudy (anonymous):

The formula for the surface area of a cylinder is A=2pr(r+h).Find h if A=80p and r=5.

OpenStudy (anonymous):

A=2pr(r+h) \[{80\pi \div(2\times5\times \pi)}-5=h, (80\div10)-5=h, 8-5=h,3=h\]

OpenStudy (anonymous):

got it grl??